OpenStudy (caozeyuan):

The roots of the equation x^3+px^2+qx+r=0 are kb, b, b/k. where all alphabet, except x, are non-zero real constants. Prove that b=-q/p, hence deduce that rp^3=q^3

OpenStudy (caozeyuan):

some one help me please!

OpenStudy (loser66):

\[\huge x^3+Px^2 +Qx +R=0~and~kb,~b,~\frac{b}{k}~are~roots\\therefore\] \[\huge kb+b+\frac{b}{k}=P\\\huge kb*b+kb*\frac{b}{k}+b*\frac{b}{k}=Q\\\huge kb*b*\frac{b}{k}=R\]

OpenStudy (loser66):

from the first two, you solve for b, and then combine with the last one to get the answer for the last part. Hope this help

OpenStudy (caozeyuan):

thanks!!!!! you saved my life!

OpenStudy (loser66):

yw